The lifetime of a machine is continuous on the interval (0, 40) with probability density function f, where f(t) is proportional to (t + 10)−2 , and t is the lifetime in years. Calculate the probability that the lifetime of the machine part is less than 10 years. Hint: Show that f(t) is legitimate and find the proportionality constant.
1
Expert's answer
2020-05-12T18:37:54-0400

"\text {Let k is a constant, then}\\

f(t)=k[(t+10)-2]=k(t+8), 0<t<40\\

k\int _0^{40}(t+8)dt=1\\

k\left[\frac{t^2}{2}+8t\right]^{40}_0=1\\

k[\frac{40^2}{2}+8(40)]=1\\

k(1120)=1\\

\therefore k=\frac{1}{1120}\\

P(t<10)=\frac{1}{1120}\int _0^{10}(t+8)dt\\

=\frac{1}{1120}\left[\frac{t^2}{2}+8t\right]^{10}_0\\

=\frac{1}{1120}[\frac{10^2}{2}+8(10)]\\

=\frac{1}{1120}\times 130\\

=\frac{13}{112}"